-3x + (D+2)y=e-t

1
Expert's answer
2022-04-05T13:29:54-0400

"y'+2y=e^{-x}+3x"

The solutions to a nonhomogeneous equation are of the form

"y(x) = y_c(x) + y_p(x)" ,

where "y_c" is the general solution to the associated homogeneous equation and "y_p" is a particular solution.

The associated homogeneous equation:

"y'+2y=0"

The general solution of this equation is determined by the roots of the characteristic equation:

"\lambda+2=0"

"\lambda=-2"

"y_c(x)=Ce^{-2x}" .

The particular solution of the differential equation:

"y_p(x)=Ae^{-x}+(Bx+C)"

"y_p'(x)=-Ae^{-x}+B"

Now put these into the original differential equation to get:

"-Ae^{-x}+B+2(Ae^{-x}+(Bx+C))=e^{-x}+3x"

Equating coefficients, we get

"A=1";   "B=\frac32"; "C=-\frac{B}{2}=-\frac34".

"y(x)=y_c(x)+y_p(x)=Ce^{-2x}+e^{-x}+\frac32x-\frac34"

Answer: "y(x)=Ce^{-2x}+e^{-x}+\frac32x-\frac34".